PartnerObserving Log for 2007 Feb 18/19

Observer1, Nicolas Bouche, Peter Buschkamp Observer2, Paul Smith

Inst. Support, Dave Thompson TO; Aaron Ceranski

UT 03:00 still waiting for the clouds to clear up;

-- NicolasBouche - 19 Feb 2007

2100: LBT open

2110-2145: Various minor issues getting wavefront sensing (WF) to converge (e.g., some manual spherical needed to be added so that software can do the colculations).

2145: Off to WF at first science field. And start fist OB of night (fuobj20_1_g). Seeing/image quality variable at 1.3-1.8".

2225: OB=fuobj20_2_g after WF. Exp 1-4: seeing 1.2".

2255: OB=fuobj20_1_r after WF. Seeing 1.1-1.3.

2325: OB=fuobj20_2_r after WF. Lose focus on exp#2. OB aborted after 3 exposures. Focus too soft. It is felt that this OB can be redone later in the run as the moon won't affect it as much as it will for many other OBs later in the run.

0000: Try M67. OB=WDs_M67_SCI_B2. Seeing really not good enough (1.6-1.7"). Finish OB, but not sure that it will be useful.

0040: Few options in queue with seeing getting worse and heavier clouds moving in. Move north to M81 for an attempt at a relatively short OB before the clouds get here.

0050: start OB=KochanekM81V. Seeing: 1.2-1.4".

0100: Staying to the north to avoid clouds.

0110: Do a few WFs, then OB=IC2574_V2, Seeing: 1.3" at start. Clouds moving in! Exposures taken through some pretty thick clouds.

0130: Close for clouds.
Topic revision: r7 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback